The College Board reports that the nationwide mean SAT Math score was 515 in 2007. Assume that SAT scores are normally distributed with a standard deviation of 116.
A.) What is the probability that a randomly selected student got an SAT score less than 500.
B.) Suppose you choose samples of SAT Math scores of size 16. Find the random sample mean and the random sample standard deviation.
C.) What is the probability that a sample of 16 students will have a sample mean SAT Math score below 500?

Answer by stanbon(75887)

The College Board reports that the nationwide mean SAT Math score was 515 in 2007. Assume that SAT scores are normally distributed with a standard deviation of 116.
A.) What is the probability that a randomly selected student got an SAT score less than 500.
z(500) = (500-515)/116 = -0.1293
P(x < 500) = P(z < -0.1293) = 0.4486
------------------------------------------------
B.) Suppose you choose samples of SAT Math scores of size 16. Find the random sample mean and the random sample standard deviation.
C.) What is the probability that a sample of 16 students will have a sample mean SAT Math score below 500?
z(500) = (500-515)/(116/sqrt(15)) = -0.5172
---
P(x-bar < 500) = P(z < -0.5172) = 0.3025
============================================
Cheers,
Stan H.
=================